# TRec: Egocentric Action Recognition using 2D Point Tracks

### v1

@@ 1 Introduction
The task of action recognition involves identifying and classifying human actions from video data. In the egocentric setting, videos are captured from a first-person perspective, typically using head-mounted cameras. Unlike exocentric (third-person) videos, which offer a stable, external view of actors performing generic activities such as sports or daily tasks, egocentric videos focus on fine-grained, domain-specific actions centered around hand–object interactions and the wearer’s visual attention. This perspective introduces unique challenges, including strong camera motion, frequent occlusions, and limited field of view, all of which complicate the modeling of both actor and object dynamics.
@@
  • We demonstrate that these tracks capture meaningful motion patterns such as hand, object, and head movements without requiring explicit semantic detection.
@@ 2 Related Work
RGB-based approaches [15, 1, 9, 5, 16, 19] rely solely on raw visual appearance for action recognition, without incorporating explicit motion or geometric cues. Many of these high-performing models are trained on large-scale datasets, often leveraging extensive pretraining and carefully designed architectures. In contrast, we intentionally adopt a simple model design to clearly demonstrate the impact of incorporating point tracks. Girdhar et al. [5] propose the Anticipative Video Transformer, which models temporal dependencies and anticipates future actions in egocentric video sequences. Their results show that the model implicitly attends to image regions containing hands and objects, even without explicit supervision for their detection. Tong et al. [16] employ masked autoencoder pretraining to improve video representation learning, achieving strong performance on egocentric and other video benchmarks. Wang et al. [19] further extend pretrained video backbone models through an interactive prototype learning framework, enhancing active-object representations by leveraging actor motion cues within RGB inputs.
RGB-based approaches [15, 1, 9, 5, 16, 19] rely solely on raw visual appearance for action recognition, without incorporating explicit motion or geometric cues. Many of these high-performing models are trained on large-scale datasets, often leveraging extensive pretraining and carefully designed architectures. In contrast, we deliberately adopt a simple model design to isolate and clearly demonstrate the effectiveness of incorporating 2D point tracks. Unlike RGB-based features that primarily encode texture and color statistics, point trajectories provide explicit geometric and temporal information, capturing object and camera motion directly from the image sequence.
Girdhar et al. [5] propose the Anticipative Video Transformer, which models temporal dependencies in egocentric videos to predict upcoming actions. Their model implicitly attends to image regions containing hands and objects, and its performance improves with increased temporal context. However, the paper does not clarify whether motion is inferred implicitly through appearance changes between frames. In contrast, our approach introduces motion information explicitly by integrating 2D point tracks, enabling the model to reason directly about the displacement and dynamics of scene elements rather than relying on learned appearance differences. This explicit representation of motion complements the visual features and enhances interpretability in modeling fine-grained egocentric actions.
Tong et al. [16] employ masked autoencoder pretraining to improve video representation learning, achieving strong performance on egocentric and other video benchmarks. Wang et al. [19] further extend pretrained video backbone models through an interactive prototype learning framework, enhancing active-object representations by leveraging actor motion cues within RGB inputs.
@@ 3 Action Recognition using Point Tracks
We present TRec, a transformer-based model for egocentric action recognition that jointly leverages 2D point tracks and image representations. TRec tracks both background and foreground points throughout a video to capture rich motion cues without relying on explicit object or hand detections. The model employs a lightweight image encoder to extract spatial features from individual frames and a transformer architecture [17] to integrate temporal dependencies across frames. A set of randomly sampled 2D points is initialized within each video, covering both foreground and background regions, and their trajectories are tracked over time. The transformer receives as input the encoded image features together with these 2D point tracks, enabling it to reason jointly about appearance and motion information. A class token and multi-headed attention pooling are used to aggregate the transformer outputs, and the resulting representation is passed through an MLP-based action recognition head to predict the action depicted in the video.
We evaluate TRec on the Something-Something-v2 [6] dataset.
@@ 3.1 Dataset Curation and Preprocessing
Something-Something-v2. The Something-Something-v2 dataset [6] contains short video clips depicting human-object interactions recorded in a wide variety of everyday environments. It comprises 174 action templates describing generic activities of “someone doing something,” such as pushing, pulling, or placing objects. It offers greater diversity in both visual context and action types, as well as a substantially large number of annotated samples, making it a valuable benchmark for training data-driven models. Because the camera remains largely stationary across most recordings, the dataset provides a stable visual setting that facilitates the analysis of object and hand dynamics without the confounding effects of egocentric motion. We use the dataset in its original form without any additional preprocessing or modifications.
Point Track Generation. We curate 2D point trajectories using CoTracker3 [7]. For each action sequence, we randomly sample 900 points in a covering the entire image, including both foreground and background regions. Video frames are sampled at 30 fps to obtain smooth trajectories. The resulting point coordinates are normalized following the procedure described in [2]. For each video we track a set of 900 points $p^{i}=(x^{i},y^{i})\in\mathbf{R}^{2}$. Here, $i\in[0,P]$ denotes the $i^{\text{th}}$ point with $P$ being the total number of points. The point tracks are denoted as $p_{t}^{i}\in\mathbf{R}^{T\times N\times 2}$ with $t\in[0,T]$ refering to the frame index, with $T$ denoting the number of frames in the video.
@@ 3.2 Network Architecture
An overview of the proposed network architecture is shown in Figure 2. The input to the network is a video $v\in\mathbf{R}^{T\times H\times W\times 3}$ where $H,W$ correspond to the image height and width, respectively. Additionally, the model receives 2D point tracks $p^{i}_{t}$ extracted from the video. The architecture consists of an image encoder $I(v_{t})$ with $v_{t}\in\mathbf{R}^{h\times w\times 3}$ as backbone and a transformer module $G:(i,p)\rightarrow g$ which takes the image features $i$ extracted by the encoder and the point tracks $t$ as input. The output $g$ of the transformer is pooled by an additional multi-head attention layer $M:g\rightarrow o$ which aggregates information from all tokens. Finally, the class token is passed through an MLP-based classification head to predict the action label.
@@ 3.3 Action Recognition
Within the transformer, spatio-temporal dependencies are modeled through self-attention across both frame and point dimensions. This mechanism allows the network to focus on motion patterns that are most informative for distinguishing between actions such as subtle object deformations or coordinated hand movements without requiring explicit hand or object detection. The multi-head attention pooling layer then aggregates the sequence of token embeddings into a single representation guided by the class token. The resulting feature vector encodes both global context and localized motion cues relevant to the observed activity.
@@ 4 Experiments
We conduct a series of experiments to evaluate the effectiveness of the proposed model under various configurations. All models are implemented using a ResNet18 backbone combined with a Transformer architecture. The Transformer module consists of six attention layers with four heads each and an intermediate dimension of 2048. Before being fed into the Transformer, the 2D point tracks $p$ are normalized and reshaped to $(B,T\times 2,P)$, after which they are passed through an MLP to match the input dimensionality of the Transformer. The number of frames per video is set to 8 if not stated otherwise. The number of point tracks is randomized between 200 and 400 points for each mini-batch.
We compare two model variants: (1) a baseline using only RGB input, (2) a trackpoint model that incorporates 2D point trajectories, and
The baseline receives only the video frames as input. Unless otherwise stated, training is performed with a batch size of 32, AdamW optimizer [10], a learning rate of $1\times 10^{-4}$ for the Transformer, and $1\times 10^{-5}$ for the backbone network. We use a cosine warm-restart learning rate scheduler with 5000 warm-up iterations and train for 20 epochs on an NVIDIA A40 GPU. Input frames are resized to $256\times 256$ and augmented using random cropping, horizontal flipping, Gaussian blur, and color jitter. The tracks are agumented according to the augmentation of their corresponding image. The observation horizon is set to $8$ frames. Out of the 900 tracked points generated per video, 400 points are randomly selected and provided to the network to promote diversity. The random number generators are seeded in all tests, thus in each experiment the model receives the same point tracks.
@@ 4.1 Track vs no Track
The first experiment compares the performance of TRec with a baseline model that shares the same architecture and training parameters. The results, summarized in Table 1, demonstrate that TRec achieves substantially higher accuracy than the baseline. This improvement indicates that the model effectively leverages the 2D point tracks to capture motion dynamics and accurately infer the actions occurring in the video.
@@
TRec $61.10\%\pm 8.66\%$ $83.95\%\pm 6.62\%$
baseline $30.27\%\pm 8.05\%$ $53.24\%\pm 8.75\%$
These models are trained with 8 frames with their corresponding point tracks as input.
@@ 4.2 Number of points
Table 2: Ablation on the number of tracked points. Performance is reported on the Something-Something-v2 datasets.
@@
4.3 Background Motion Filtering Using Kernel Density Estimation
@@
vanilla TRec $61.10\%$ $83.95\%$
vanilla TRec + KDE $52.46\%$% $74.04\%$
FilterTRec + KDE $36.34\%$ $60.57\%$
@@ 4.4 Single-Image Evaluation
In this setup, only the initial image of each video is provided as input during training and testing. Because this frame may not contain visible hands or objects, the model must rely entirely on the tracked motion to infer hand movements and object interactions. Moreover, the Something-Something-v2 dataset includes several action categories that can only be recognized by observing the final frames of a video, such as pretending to do something or revealing something. Since the model cannot predict these actions without the last frame, we evaluate the approach both on the unmodified dataset and on a version with these ambiguous categories removed, specifically those containing the terms pretend, uncover, reveal, fail in their annotations.
@@ 5 Discussion
The experimental results demonstrate that incorporating 2D point tracks provides a clear advantage for egocentric action recognition compared to purely RGB-based methods. By explicitly encoding motion cues through tracked points, our model captures spatial-temporal dynamics that RGB alone cannot represent. This finding supports our initial hypothesis that motion trajectories offer a lightweight and interpretable alternative to computationally demanding approaches such as optical flow or hand-object detection.
@@
The single-image experiment further highlights the discriminative power of the point trajectories. Even when only the first frame of a video is available, the model achieves higher accuracy than the RGB-only baseline, demonstrating that temporal motion encoded by the tracks carries sufficient information to infer the underlying action. This result also underlines the interpretability of our representation where the model can reason about hand and object movement even if they are not visible in the initial frame.
Finally, the kernel density filtering experiment reveals that background motion contributes significantly to action recognition in egocentric videos. Removing background points leads to a noticeable drop in accuracy, indicating that the model uses global motion patterns such as camera or head movements as contextual signals for understanding actions. This aligns with the egocentric nature of the data, where background dynamics often correlate with the actor’s motion and intention.

### v2

@@ 1 Introduction
The task of action recognition involves identifying and classifying human actions from video data. In the egocentric setting, videos are captured from a first-person perspective, typically using head-mounted cameras. Unlike exocentric (third-person) videos, which offer a stable, external view of actors performing generic activities such as sports or daily tasks, egocentric videos focus on fine-grained, domain-specific actions centered around hand-object interactions and the wearer’s visual attention. This perspective introduces unique challenges, including strong camera motion, frequent occlusions, and limited field of view, all of which complicate the modeling of both actor and object dynamics.
@@
  • We demonstrate that these tracks capture meaningful motion patterns spanning hand, object, and head movements, without requiring explicit semantic detection.
@@ 2 Related Work
RGB-based approaches [15, 1, 9, 5, 16, 19] rely solely on raw visual appearance for action recognition, without incorporating explicit motion or geometric cues. Many of these high-performing models are trained on large-scale datasets, often leveraging extensive pretraining and carefully designed architectures. In contrast, we deliberately adopt a simple model design to isolate and clearly demonstrate the effectiveness of incorporating 2D point tracks. Unlike RGB-based features that primarily encode texture and color statistics, point trajectories provide explicit geometric and temporal information, capturing object and camera motion directly from the image sequence. Girdhar et al. [5] propose the Anticipative Video Transformer, which models temporal dependencies in egocentric videos to predict upcoming actions. Their model implicitly attends to image regions containing hands and objects, and its performance improves with increased temporal context. However, the paper does not clarify whether motion is inferred implicitly through appearance changes between frames. In contrast, our approach introduces motion information explicitly by integrating 2D point tracks, enabling the model to reason directly about the displacement and dynamics of scene elements rather than relying on learned appearance differences. This explicit representation of motion complements the visual features and enhances interpretability in modeling fine-grained egocentric actions. Tong et al. [16] employ masked autoencoder pretraining to improve video representation learning, achieving strong performance on egocentric and other video benchmarks. Wang et al. [19] further extend pretrained video backbone models through an interactive prototype learning framework, enhancing active-object representations by leveraging actor motion cues within RGB inputs.
@@ 3 Action Recognition using Point Tracks
We present TRec, a transformer-based model for egocentric action recognition that jointly leverages 2D point tracks and image representations. TRec tracks both background and foreground points throughout a video to capture rich motion cues without relying on explicit object or hand detections. The model employs a lightweight image encoder to extract spatial features from individual frames and a transformer architecture [17] to integrate temporal dependencies across frames. A set of randomly sampled 2D points is initialized within each video, covering both foreground and background regions, and their trajectories are tracked over time. The transformer receives as input the encoded image features together with these 2D point tracks, enabling it to reason jointly about appearance and motion information. An additional multi-headed attention layer is used to aggregate the transformer outputs, and the resulting representation is passed through an MLP-based action recognition head to predict the action depicted in the video. We evaluate TRec on the Something-Something-v2 [6] dataset.
@@ 3.1 Dataset Curation and Preprocessing
Something-Something-v2. The Something-Something-v2 dataset [6] contains short video clips depicting human-object interactions recorded in a wide variety of everyday environments. It comprises 174 action templates describing generic activities of someone doing ”something”, such as pushing, pulling, or placing objects. It offers greater diversity in both visual context and action types, as well as a substantially large number of annotated samples, making it a valuable benchmark for training data-driven models. Because the camera remains largely stationary across most recordings, the dataset provides a stable visual setting that facilitates the analysis of object and hand dynamics without the confounding effects of egocentric motion. We use the dataset in its original form without any additional preprocessing or modifications.
Point Track Generation. We curate 2D point trajectories using CoTracker3 [7]. For each action sequence, we randomly sample 900 points covering the entire image, including both foreground and background regions. Video frames are sampled at 30 fps to obtain smooth trajectories. The resulting point coordinates are normalized following the procedure described in [2]. For each video we track a set of 900 points $p^{i}=\{(x^{i}_{t},y^{i}_{t})\in\mathbb{R}^{2}|t=0,...T\}$. Here, $i\in[0,P]$ denotes the $i^{\text{th}}$ point with $P$ being the total number of points while $t\in[0,T]$ refrs to the frame index, with $T$ denoting the number of frames in the video. We refer to a point at time step $t$ as $p_{t}^{i}\in\mathbb{R}^{2}$. The set of all point tracks within a video is denoted as $\mathbf{P}$.
@@ 3.2 Network Architecture
An overview of the proposed network architecture is shown in Figure 2. The input to the network is a video $v\in\mathbb{R}^{T\times H\times W\times 3}$ where $H,W$ correspond to the image height and width, respectively. Additionally, the model receives 2D point tracks $\mathbf{P}$ extracted from the video. The architecture consists of an image encoder $I(v_{t})$ with $v_{t}\in\mathbb{R}^{h\times w\times 3}$ as backbone and a transformer module $G:(i,p)\rightarrow g$ which takes the image features $i$ extracted by the encoder and the point tracks $t$ as input. The output $g$ of the transformer is pooled by an additional multi-head attention layer $M:g\rightarrow o$ which aggregates information from all tokens. Finally, the class token is passed through an MLP-based classification head to predict the action label.
@@ 3.3 Action Recognition
Within the transformer, spatio-temporal dependencies are modeled through self-attention across both frame and point dimensions. This mechanism allows the network to emphasize motion patterns most relevant for action discrimination, encompassing subtle object deformations and coordinated hand movements, without requiring explicit hand or object detection. The multi-head attention pooling layer then aggregates the sequence of token embeddings into a single representation guided by the class token. The resulting feature vector encodes both global context and localized motion cues relevant to the observed activity.
@@ 4 Experiments
We conduct a series of experiments to evaluate the effectiveness of the proposed model under various configurations. All models are implemented using a ResNet18 backbone combined with a Transformer architecture. The Transformer module consists of six attention layers with four heads each and an intermediate dimension of 2048. Before being fed into the Transformer, the 2D point tracks $p$ are normalized and reshaped to $(B,T\times 2,P)$, then passed through an MLP to project them to the Transformer’s input dimensionality. Here, $B$ denotes the batch size. The number of frames per video is set to 8 if not stated otherwise. The number of point tracks is randomized between 200 and 400 points for each mini-batch.
We compare two model variants: (1) a baseline using only RGB input, (2) a trackpoint model that incorporates 2D point trajectories
The baseline receives only the video frames as input. Unless otherwise stated, training is performed with a batch size of 32, AdamW optimizer [10], a learning rate of $1\times 10^{-4}$ for the Transformer, and $1\times 10^{-5}$ for the backbone network. We use a cosine warm-restart learning rate scheduler with 5000 warm-up iterations and train for 20 epochs on an NVIDIA A40 GPU. Input frames are resized to $256\times 256$ and augmented using random cropping, horizontal flipping, Gaussian blur, and color jitter. The tracks are agumented according to the augmentation of their corresponding image. The observation horizon is set to $8$ frames. Out of the 900 tracked points generated per video, 400 points are randomly selected and provided to the network to promote diversity. The random number generators are seeded in all tests, thus in each experiment the tested model receives the same point tracks.
@@ 4.1 Track vs no Track
The first experiment compares the performance of TRec with a baseline model that shares the same architecture and training parameters but only receives the RGB frames as input. The results, summarized in Table 1, demonstrate that TRec achieves substantially higher accuracy than the baseline. This improvement indicates that the model effectively leverages the 2D point tracks to capture motion dynamics and accurately infer the actions occurring in the video.
@@
TRec $61.10\pm 8.66$ $83.95\pm 6.62$
baseline $30.27\pm 8.05$ $53.24\pm 8.75$
@@ 4.2 Number of points
Table 2: Ablation on the number of tracked points
@@
4.3 Background Motion as Context for Action Recognition
@@
vanilla TRec $61.10$ $83.95$
vanilla TRec + KDE $52.46$ $74.04$
FilterTRec + KDE $36.34$ $60.57$
@@ 4.4 Single-Image Evaluation
In this setup, only the initial image of each video is provided as input during training and testing. Because this frame may not contain visible hands or objects, the model must rely entirely on the tracked motion to infer hand movements and object interactions. Moreover, the Something-Something-v2 dataset includes several action categories that can only be recognized by observing the final frames of a video, such as pretending to do something or revealing something. As the model cannot predict these actions without the final frame, we evaluate the approach on the original dataset and on a version excluding ambiguous categories whose annotations contain pretend, uncover, reveal, or fail.
@@ 5 Discussion
The experimental results demonstrate that incorporating 2D point tracks provides a clear advantage for egocentric action recognition compared to purely RGB-based methods. By explicitly encoding motion cues through tracked points, our model captures spatial-temporal dynamics that RGB alone cannot represent. This finding supports our initial hypothesis that motion trajectories offer a lightweight and interpretable alternative to computationally demanding approaches such as optical flow or hand–object detection.
@@
The single-image experiment further highlights the discriminative power of the point trajectories. Even when only the first frame of a video is available, the model achieves higher accuracy than the RGB-only baseline, demonstrating that temporal motion encoded by the tracks carries sufficient information to infer the underlying action. This result also underlines the interpretability of our representation, in which the model can reason about hand and object movement even when they are not visible in the initial frame.
Finally, the kernel density filtering experiment reveals that background motion contributes significantly to action recognition in egocentric videos. Removing background points leads to a noticeable drop in accuracy, suggesting that the model relies on global motion patterns, encompassing camera and head movements, as contextual cues for action understanding. This aligns with the egocentric nature of the data, where background dynamics often correlate with the actor’s motion and intention.